# LiDAR Point Cloud-based Multiple Vehicle Tracking with Probabilistic Measurement-Region Association

### v2

@@ III-D PMRA-PMBM Update
In the PMRA-PMBM update procedure, given the predicted density of a target $\mathcal{F}(\mathbf{x}_{+};\alpha_{+},\beta_{+},\Xi_{+})$ and a set of measurements $\mathbf{Y}$ associating with the target, the updated density $\mathcal{F}(\mathbf{x};\alpha,\beta,\Xi)$ is obtained as follows (the time index $k$ is omitted). First, the gamma parameters of measurement rate are updated with the Bayesian recursion in [18]:
@@
where $\Gamma(\cdot)$ is the gamma function, $\mathcal{L}_{\gamma}$ is the predictive likelihood. Then, the updated particles $\Xi$ are obtained from a simplified PMRA model, which processes the measurements in $\mathbf{Y}$ one by one instead of in a batch. In this way, the total number of measurement-region associations is reduced from $a=5^{|\mathbf{Y}|}$ to $a=5{|\mathbf{Y}|}$. Assign an arbitrary index order to the measurement set, i.e., $\mathbf{Y}=\{\mathbf{z}^{m}\}_{m=1}^{|\mathbf{Y}|}$. Then, for the measurement $\mathbf{z}^{m}$, the particle weights are updated by
@@
where $\Xi_{-}=\{\mathbf{m}_{-}^{(l)},E_{-}^{(l)},w_{-}^{(l)}\}_{l=1}^{L}$ represents the particles and weights calculated with $\mathbf{z}^{m-1}$, and $\Xi_{-}=\Xi_{+}$ for $m=1$. The weights are normalized such that $\sum_{l}w^{(l)}=1$. Here, the conditional measurement likelihoods $\mathcal{L}_{r_{n}}$ are defined by (II-C) and (16); the prior association probability $P\{r_{n}|\mathbf{m}_{-}^{(l)},E_{-}^{(l)}\}$ is calculated with (II-D). If the effective particle number $1/\sum_{l}(w^{(l)})^{2}$ is below a threshold $L_{\mathrm{e}}$, systematic resampling is performed to resample the particles and avoid the particle degeneracy [25]. After all the measurements are processed, the predictive likelihood of particles is approximated by
@@
In stage one, the algorithm performs clustering and gating with the predicted PMBM density and the measurement set. The clusters $\{\mathbf{Y}^{b,h}\}_{h\in\mathbf{H}^{b}}$ considered as possibly originating from the newborn targets are determined with the method in Section III-A. Note that the gating threshold $d_{\mathrm{out}}$ should be decreased if the target birth frequently occurs close to existing targets, and $d_{\mathrm{in}}$ should be increased for maneuvering targets.
In stage two, the algorithm calculates the density for undetected targets by updating the predicted PPP intensity $D_{+}^{u}$. Since the target is not associated with any measurement, the particles are directly propagated, and the mixture reduction method in [18] is used to calculate the gamma parameters.
In stage three, the algorithm generates data association hypotheses based on the clustering and gating results, then update the MBM density of the detected targets. Under the $j_{+}$-th global hypothesis, a cluster $\mathbf{Y_{S}}$ not associating with any existing targets (i.e, $\mathbf{S}\cap\mathbf{I}^{j_{+}}=\emptyset$) is used to calculate the density of a target detected for the first time. The updated density is multi-modal with one mode for each of the Poisson components in $D_{+}^{u}$. Therefore, gamma mixture reduction and
@@
Input: $D_{+}^{u},\{(C_{+}^{j},\{(r_{+}^{j,i},f_{+}^{j,i})\}_{i\in\mathbf{I}_{+}^{j}})\}_{j\in\mathbf{J}_{+}}$, measurement set $\mathbf{Z}$.
Output: $D^{u},\{(C^{j},\{(r^{j,i},f^{j,i})\}_{i\in\mathbf{I}^{j}})\}_{j\in\mathbf{J}}$.
From $D_{+}^{u},\{(C_{+}^{j},\{(r_{+}^{j,i},f_{+}^{j,i})\}_{i\in\mathbf{I}^{j}})\}_{j\in\mathbf{J}}$ and $\mathbf{Z}$, compute clustering and gating results $\mathcal{CG}$.
Initialize: $D^{u}\leftarrow 0$, $D^{b}\leftarrow 0$, $\mathbf{J}\leftarrow\emptyset$, $j\leftarrow 0$, $h\leftarrow 0$.
for $h\in\mathbf{H}_{+}^{u}$ do: [PPP update for miss detected targets]
$q_{1}=1-p_{\mathrm{D}}(\hat{\mathbf{x}}_{+}^{u,h})$, $q_{2}=p_{\mathrm{D}}(\hat{\mathbf{x}}_{+}^{u,h})[\beta_{+}^{u,h}(\beta_{+}^{u,h}+1)]^{\alpha_{+}^{u,h}}$,
$q_{\mathrm{D}}=q_{1}+q_{2}$, $\beta^{u,h}=1/\{q_{1}/(q_{\mathrm{D}}\beta_{+}^{u,h})+q_{2}/[q_{\mathrm{D}}(\beta_{+}^{u,h}+1)]\}$,
$w^{u,h}=q_{\mathrm{D}}w_{+}^{u,h}$. $D^{u}\leftarrow D^{u}+w^{u,h}\mathcal{F}(\alpha_{+}^{u,h},\beta^{u,h},\Xi_{+}^{u,h})$.
@@
 Compute associations $\mathbb{A}^{j_{+}}$ based on $\mathcal{CG}$.
@@
    for $h\in\mathbf{H}^{u}_{+}$ do:
     From $\alpha^{u,h}_{+},\beta^{u,h}_{+},\Xi^{u,h}_{+}$ and $\mathbf{S}$, compute $\alpha^{u,h}$, $\beta^{u,h}$, $\Xi^{u,h}$,
     $\mathcal{L}_{\gamma}^{u,h}$, $\mathcal{L}_{\Xi}^{u,h}$ as in (28), (29), (30).
@@
    From $\{\alpha^{u,h},\beta^{u,h}\}_{h\in\mathbf{H}^{u}_{+}}$, compute $\alpha$, $\beta$ with gamma mix-
    ture reduction.
    From $\{\Xi^{u,h}\}_{h\in\mathbf{H}^{u}_{+}}$, compute $\Xi$ with systematic resampling.
@@
    $\mathcal{L}=\sum_{h\in\mathbf{H}^{u}_{+}}w^{u,h}p_{\mathrm{D}}(\hat{\mathbf{x}}^{u,h})\mathcal{L}_{\gamma}^{u,h}\mathcal{L}_{\Xi}^{u,h}$.
@@
From $\mathcal{CG}$, compute clusters $\{\mathbf{Y}^{b,h}\}_{h\in\mathbf{H}^{b}}$.
for $h\in\mathbf{H}^{b}$ do: [PPP initialization for new birth targets]
 Initialize: From $\mathbf{Y}^{b,h}$, compute $\Xi^{b,h}$ by (III-B), (III-B), and (25).
 Increment: $D^{b}\leftarrow D^{b}+w^{b}\mathcal{F}(\mathbf{x};\alpha^{b},\beta^{b},\Xi^{b,h})$.
@@
$D^{u}\leftarrow D^{u}+D^{b}$, where $\mathbf{H}^{u}\leftarrow\mathbf{H}_{+}^{u}\cup\mathbf{H}^{b}$.
@@
In the final stage, the algorithm initializes the birth PPP intensity $D^{b}$ with the clusters $\{\mathbf{Y}^{b,h}\}_{h\in\mathbf{H}^{b}}$ and combines it into the updated PPP density $D^{u}$.
@@ IV Simulation Study
In this section, the tracking performance of PMRA-PMBM algorithm is evaluated with simulated data. As illustrated in Fig. 3, we simulate a typical application scenario of the intelligent transportation system, where a LiDAR sensor is mounted on the road side unit to monitor the traffic at the intersection [26][27]. During the $20\mathrm{s}$ simulation period, $6$ vehicles with $4.5\mathrm{m}\times 1.8\mathrm{m}$ rectangular extent enter and leave the surveillance area at different time steps. A stationary LiDAR sensor with $2\mathrm{Hz}$ sample rate and $0.5^{\circ}$ angular resolution collects measurement points of the vehicles. We assume that the target-originated measurements are reflected from the visible edges of extent rectangles, and the covariance of additive zero-mean Gaussian measurement noise is $\mathrm{diag}([(0.1^{\circ})^{2},(0.01\mathrm{m})^{2}])$. The clutters are uniformly distributed over the $100\mathrm{m}\times 100\mathrm{m}$ surveillance area with Poisson rate $\mu^{c}=20$.
@@
The proposed PMRA-PMBM filter is compared with the DRA-PMBM (an implementation of “Model-V” DRA method [14] under the PMBM framework) and the GGIW-PMBM [11] filters in the simulation scenario. All three algorithms use the CT model to track target kinematic state, and the prior target extent is set as $4\mathrm{m}\times 2\mathrm{m}$. The particle number and resample threshold of the PMRA-PMBM filter are $L=1000$ and $L_{\mathrm{e}}=100$. The other parameters are fine-tuned for each algorithm to obtain the optimal performance.
The GOSPA results averaged over 100 Monte Carlo runs are shown in Fig. 4 and Table IV. The PMRA-PMBM has overall lower estimation error for target position and extent than the other algorithms, especially the GGIW-PMBM which assumes measurements are Gaussian distributed over the target extent. Compared to the DRA-PMBM, the GOSPA of PMRA-PMBM decreases faster after $2$ targets appear at $5\mathrm{s}$ and exhibits less fluctuation when target cardinality changes frequently between $9\mathrm{s}$ and $15\mathrm{s}$. This indicates that the PMRA-PMBM has superior estimation accuracy and stability in a rapidly changing MTT scenario. As illustrated by Fig. 5, the PMRA-PMBM filter can accurately estimates the vehicle extent from the simulated LiDAR point cloud. However, the particle-based implementation of the PMRA-PMBM requires more compuational resources than the other two algorithms. As shown in Table IV, the average processed frames per second (FPS) of PMRA-PMBM is about 47.5% that of DRA-PMBM and 15.1% that of GGIW-PMBM on the same simulation platform.
@@ V Conclusion
This paper presents the PMRA-PMBM filter for tracking multiple vehicles with LiDAR point clouds. The proposed PMRA model improves the estimation accuracy and stability of extended target state compared with the existing DRA methods. Simulation results show that the particle-based implementation of the PMRA-PMBM filter can achieve superior estimation accuracy in both the position and extent of vehicle compared to the GGIW-PMBM and DRA-PMBM filters. In our future work, message passing methods and the parallelized particle filter will be investigated to reduce the computational complexity of the proposed algorithm.

### v3

@@ III-D PMRA-PMBM Update
In the PMRA-PMBM update procedure, given the predicted density of a target $\mathcal{F}(\mathbf{x}_{+};\alpha_{+},\beta_{+},\Xi_{+})$ and a set of measurements $\mathbf{Y}$ associated with the target, the updated density $\mathcal{F}(\mathbf{x};\alpha,\beta,\Xi)$ is obtained as follows (the time index $k$ is omitted). First, the gamma parameters of the measurement rate are updated with the Bayesian recursion in [18]:
@@
where $\Gamma(\cdot)$ is the gamma function, $\mathcal{L}_{\gamma}$ is the predictive likelihood. Then, the updated particles $\Xi$ are obtained from a simplified PMRA model, which sequentially processes the measurements $\mathbf{Y}$ in order to reduce the total number of measurement-region associations from $a=5^{|\mathbf{Y}|}$ to $a=5{|\mathbf{Y}|}$. Assign an arbitrary index order to the measurement set, i.e., $\mathbf{Y}=\{\mathbf{z}^{m}\}_{m=1}^{|\mathbf{Y}|}$. Then, for the measurement $\mathbf{z}^{m}$, the particle weights are updated by
@@
where $\Xi_{-}=\{\mathbf{m}_{-}^{(l)},E_{-}^{(l)},w_{-}^{(l)}\}_{l=1}^{L}$ represents the particles and weights calculated with $\mathbf{z}^{m-1}$, and $\Xi_{-}=\Xi_{+}$ for $m=1$. The weights are normalized such that $\sum_{l}w^{(l)}=1$. Here, the conditional measurement likelihoods $\mathcal{L}_{r_{n}}$ are defined by (15) and (16); the prior association probability $P\{r_{n}|\mathbf{m}_{-}^{(l)},E_{-}^{(l)}\}$ is calculated with (22). If the effective particle number $1/\sum_{l}(w^{(l)})^{2}$ is below a threshold $L_{\mathrm{e}}$, systematic resampling is performed to resample the particles and avoid the particle degeneracy [25]. After all the measurements are processed, the predictive likelihood of particles is approximated by
@@
In stage one, the algorithm performs clustering and gating with the predicted PMBM density and the measurement set. The clusters $\{\mathbf{Y}^{\mathrm{b},h}\}_{h\in\mathbf{H}^{\mathrm{b}}}$ considered as possibly originating from the newborn targets are determined with the method in Section III-A. Note that the gating threshold $d_{\mathrm{out}}$ should be decreased if the target birth frequently occurs close to existing targets, and $d_{\mathrm{in}}$ should be increased for maneuvering targets.
In stage two, the algorithm calculates the density for undetected targets by updating the predicted PPP intensity $D_{+}^{\mathrm{u}}$. Since the target is not associated with any measurement, the particles are directly propagated, and the mixture reduction method in [18] is used to calculate the gamma parameters.
In stage three, the algorithm generates data association hypotheses based on the clustering and gating results, then updates the MBM density of the detected targets. Under the $j_{+}$-th global hypothesis, a cluster $\mathbf{Y_{S}}$ not associated with any existing targets (i.e, $\mathbf{S}\cap\mathbf{I}^{j_{+}}=\emptyset$) is used to calculate the density of a target detected for the first time. The updated density is multi-modal with one mode for each of the Poisson components in $D_{+}^{\mathrm{u}}$. Therefore, gamma mixture reduction and
@@
Input: $D_{+}^{\mathrm{u}},\{(C_{+}^{j},\{(r_{+}^{j,i},f_{+}^{j,i})\}_{i\in\mathbf{I}_{+}^{j}})\}_{j\in\mathbf{J}_{+}}$, measurement set $\mathbf{Z}$.
Output: $D^{\mathrm{u}},\{(C^{j},\{(r^{j,i},f^{j,i})\}_{i\in\mathbf{I}^{j}})\}_{j\in\mathbf{J}}$.
From $D_{+}^{\mathrm{u}},\{(C_{+}^{j},\{(r_{+}^{j,i},f_{+}^{j,i})\}_{i\in\mathbf{I}^{j}})\}_{j\in\mathbf{J}}$ and $\mathbf{Z}$, compute clustering and gating results $\mathcal{CG}$.
Initialize: $D^{\mathrm{u}}\leftarrow 0$, $D^{\mathrm{b}}\leftarrow 0$, $\mathbf{J}\leftarrow\emptyset$, $j\leftarrow 0$, $h\leftarrow 0$.
for $h\in\mathbf{H}_{+}^{\mathrm{u}}$ do: [PPP update for undetected targets]
$q_{1}=1-p_{\mathrm{D}}(\hat{\mathbf{x}}_{+}^{\mathrm{u},h})$, $q_{2}=p_{\mathrm{D}}(\hat{\mathbf{x}}_{+}^{\mathrm{u},h})[\beta_{+}^{\mathrm{u},h}/(\beta_{+}^{\mathrm{u},h}+1)]^{\alpha_{+}^{\mathrm{u},h}}$,
$q_{\mathrm{D}}=q_{1}+q_{2}$, $\beta^{\mathrm{u},h}=1/\{q_{1}/(q_{\mathrm{D}}\beta_{+}^{\mathrm{u},h})+q_{2}/[q_{\mathrm{D}}(\beta_{+}^{\mathrm{u},h}+1)]\}$,
$w^{\mathrm{u},h}=q_{\mathrm{D}}w_{+}^{\mathrm{u},h}$. $D^{\mathrm{u}}\leftarrow D^{\mathrm{u}}+w^{\mathrm{u},h}\mathcal{F}(\alpha_{+}^{\mathrm{u},h},\beta^{\mathrm{u},h},\Xi_{+}^{\mathrm{u},h})$.
@@
 Compute possible associations $\mathbb{A}^{j_{+}}$ based on $\mathcal{CG}$.
@@
    for $h\in\mathbf{H}^{\mathrm{u}}_{+}$ do:
     From $\alpha^{\mathrm{u},h}_{+},\beta^{\mathrm{u},h}_{+},\Xi^{\mathrm{u},h}_{+}$ and $\mathbf{S}$, compute $\alpha^{\mathrm{u},h}$, $\beta^{\mathrm{u},h}$, $\Xi^{\mathrm{u},h}$,
     $\mathcal{L}_{\gamma}^{\mathrm{u},h}$, $\mathcal{L}_{\Xi}^{\mathrm{u},h}$ as in (28), (29), (30).
@@
    From $\{\alpha^{\mathrm{u},h},\beta^{\mathrm{u},h}\}_{h\in\mathbf{H}^{\mathrm{u}}_{+}}$, compute $\alpha$, $\beta$ with gamma mix-
    ture reduction [18].
    From $\{\Xi^{\mathrm{u},h}\}_{h\in\mathbf{H}^{\mathrm{u}}_{+}}$, compute $\Xi$ with systematic resampling.
@@
    $\mathcal{L}=\sum_{h\in\mathbf{H}^{\mathrm{u}}_{+}}w^{\mathrm{u},h}p_{\mathrm{D}}(\hat{\mathbf{x}}^{\mathrm{u},h})\mathcal{L}_{\gamma}^{\mathrm{u},h}\mathcal{L}_{\Xi}^{\mathrm{u},h}$.
@@
From $\mathcal{CG}$, compute clusters $\{\mathbf{Y}^{\mathrm{b},h}\}_{h\in\mathbf{H}^{\mathrm{b}}}$.
for $h\in\mathbf{H}^{\mathrm{b}}$ do: [PPP initialization for newborn targets]
 Initialize: From $\mathbf{Y}^{\mathrm{b},h}$, compute $\Xi^{\mathrm{b},h}$ by (23), (24), and (25).
 Increment: $D^{\mathrm{b}}\leftarrow D^{\mathrm{b}}+w^{\mathrm{b}}\mathcal{F}(\mathbf{x};\alpha^{\mathrm{b}},\beta^{\mathrm{b}},\Xi^{\mathrm{b},h})$.
@@
$D^{\mathrm{u}}\leftarrow D^{\mathrm{u}}+D^{\mathrm{b}}$, where $\mathbf{H}^{\mathrm{u}}\leftarrow\mathbf{H}_{+}^{\mathrm{u}}\cup\mathbf{H}^{\mathrm{b}}$.
@@
In the final stage, the algorithm initializes the birth PPP intensity $D^{\mathrm{b}}$ with the clusters $\{\mathbf{Y}^{\mathrm{b},h}\}_{h\in\mathbf{H}^{\mathrm{b}}}$ and combines it into the updated PPP density $D^{\mathrm{u}}$.
@@ IV Simulation Study
In this section, the performance of the PMRA-PMBM algorithm is evaluated with simulated data. As illustrated in Fig. 3, we simulate a typical application scenario of the intelligent transportation system, where a LiDAR sensor is mounted on the road side unit to monitor the traffic at the intersection [26][27]. During the $20\mathrm{s}$ simulation period, $6$ vehicles with $4.5\mathrm{m}\times 1.8\mathrm{m}$ rectangular extent enter and leave the surveillance area at different time steps. A stationary LiDAR sensor with $2\mathrm{Hz}$ sample rate and $0.5^{\circ}$ angular resolution collects measurements of the vehicles. We assume that the target-originated measurements are reflected from the visible edges of the extent rectangles, and the covariance of the additive zero-mean Gaussian measurement noise is $\mathrm{diag}([(0.1^{\circ})^{2},(0.01\mathrm{m})^{2}])$. The clutters are uniformly distributed over the $100\mathrm{m}\times 100\mathrm{m}$ surveillance area with Poisson rate $\mu^{c}=20$.
@@
The proposed PMRA-PMBM filter is compared with the DRA-PMBM (an implementation of “Model-V” DRA method [14] under the PMBM framework) and the GGIW-PMBM [11] filters in the simulation scenario. All three algorithms use the CT model to track the kinematic state of targets, and the prior target extent is set as $4\mathrm{m}\times 2\mathrm{m}$. The particle number and resample threshold of the PMRA-PMBM filter are $L=1000$ and $L_{\mathrm{e}}=100$. The other parameters are fine-tuned for each algorithm to obtain the optimal performance.
The GOSPA results averaged over 100 Monte Carlo runs are shown in Fig. 4 and Table IV. The PMRA-PMBM has lower estimation error for positions and extents than the other two algorithms, especially the GGIW-PMBM which assumes measurements are Gaussian distributed around the target’s center. Compared to the DRA-PMBM, the GOSPA of PMRA-PMBM decreases faster after $2$ targets appear at $5\mathrm{s}$ and exhibits less fluctuation when the target cardinality changes frequently between $9\mathrm{s}$ and $15\mathrm{s}$. This indicates that the PMRA-PMBM has superior estimation accuracy and stability in a rapidly changing MTT scenario. As illustrated by Fig. 5, the PMRA-PMBM filter can accurately estimates the vehicle extent from the simulated LiDAR point cloud. However, the particle-based implementation of the PMRA-PMBM requires more compuational resources than the other two algorithms. As shown in Table IV, the average processed frames per second (FPS) of the PMRA-PMBM is about 47.5% that of the DRA-PMBM, and 15.1% that of the GGIW-PMBM on the same simulation platform.
@@ V Conclusion
This paper presents the PMRA-PMBM filter for tracking multiple vehicles with LiDAR point clouds. The proposed PMRA model improves the estimation accuracy and stability of the extended target state compared with the existing DRA methods. Simulation results show that the particle-based implementation of the PMRA-PMBM filter can achieve superior estimation accuracy in both the position and extent of vehicle compared to the GGIW-PMBM and DRA-PMBM filters. In our future work, message passing methods and the parallelized particle filter will be investigated to reduce the computational complexity of the proposed algorithm.